OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = -3x^2 + 12x – 7

OpenStudy (anonymous):

@Jack175

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@kyusakazaki

OpenStudy (anonymous):

@xo_ellaaavasquez coming

OpenStudy (anonymous):

??

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

i think -2, -43

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

thats one of my answers

OpenStudy (anonymous):

x=12/-2*-3=2 ,y=-3*4 +2*12 -7=-12+24-7=5 ,then (2,5)

OpenStudy (anonymous):

yes thats correct thanks

OpenStudy (anonymous):

@xo_ellaaavasquez , you are welcome, sorry if i was late , but the administration was messaging me about the honor code

OpenStudy (anonymous):

its ok i have another problem can you help please